# Enhancement of non-Gaussianity and nonclassicality of pair coherent states with postselected von Neumann measurement

### v1

@@ VIII Discussion
(4) Weak coupling strength parameter ($\Gamma\ll 1$) with small state parameter $\gamma$ ($|\gamma|\ll 1$). Here, as both the coupling parameter $\Gamma\ll 1$ and $\gamma$ are very small, we can neglect $\Gamma\gamma^{1+\delta/2}$ and higher-order terms in our system state. In this extreme case, the expression for our output state $|\Psi\rangle$ further changes to (unnormalized):
@@
In this regime, the WVA is dominant, and $\Gamma Re[\langle\sigma_{x}\rangle_{w}]$ can become a significant quantity compared to $\gamma$, especially if the pre and post-selected system states are nearly orthogonal. If we assume the PND equal zero ($\delta=0$), the $a$-mode is prepared in a superposition state of $|0\rangle_{a}$ and $|1\rangle_{a}$, while the $b$-mode remains in the vacuum state. Furthermore, under sufficiently large anomalous weak value, it is possible to detect a single photon from the $a$-mode with higher probability. This feature could be useful in quantum computation processes, as discussed in (Kok et al., 2007). More interestingly, in this extreme case, we can also prepare the state $\left|1+\delta\right\rangle_{a}\left|1\right\rangle_{b}$, but with a lower generation probability. For this extreme case, the advantages of our scheme are highlighted by the very small value regimes shown in the numerical results of Figs. 6(a) and 6(b).
@@
We also note that our theoretical scheme can be realized on an optical platform, as the initial system preparation, post-selection of the measured system, and weak coupling between the measured system and MD can all be accomplished in optical labs, as in previous WVA experiments (Xu et al., 2013; Zhu et al., 2021). Furthermore, it may also be possible to implement our scheme in a trapped ion system, as the von Neumann-type interaction Hamiltonian $H_{int}=g\sigma_{x}\otimes P_{x}$ can be prepared in two-dimensional harmonic oscillators of trapped ion systems (Wu et al., 2019; Pan et al., 2020; Su et al., 2023; Jeon et al., 2024; Wu et al., 2024).
@@
Finally, the postselected von Neumann measurement utilizes the technique of WVA, which enables the PCS to exhibit excellent non-classical properties. The PCS under postselected von Neumann measurement can show improved performance in squeezing, quantum statistics, and entanglement by selecting appropriate parameters. This method may provide a valuable quantum resource for enhancing the efficiency of quantum teleportation and other quantum tasks based on PCS (Tara and Agarwal, 1994; GÁBRIS and AGARWAL, 2007; Albert et al., 2019).
@@
In this paper, we have proposed a theoretical model to enhance the non-Gaussianity and nonclassicality of PCS using postselected von Neumann measurement. We apply a postselected measurement to one mode (the $a$-mode) of the PCS and investigate related properties, including squeezing, quantum statistics, and entanglement. The results show that after the postselected von Neumann measurement, the associated characteristics of the measurement output state are enhanced in the WM regime with anomalous weak values. Quadrature squeezing achieves better-squeezing effects than the PCS over a wide range of state parameters $\gamma$ under WM. The sum squeezing effect occurs due to WVA for larger weak values. Numerical results using HZ correlation and EPR correlation also show that the entanglement between the two modes of the postselected measurement-enhanced PCS becomes stronger than the initial state. We confirm these findings by examining the scaled joint Wigner function for different system parameters. We also analyze the changes in the initial state after the measurement process in terms of fidelity and observe that as the coupling strength parameter $\Gamma$ increases, the initial state $|\phi\rangle$ changes and can even become orthogonal to $|\phi\rangle$. From this perspective, we also discuss the potential applications of our scheme in quantum state preparation processes. We explored the possibility of performing postselected von Neumann measurements in multimode radiation fields. For simplicity, we only considered the postselected measurement on one mode of the PCS. It would be interesting to extend this to postselected measurements on both modes of two-mode states, such as the two-mode squeezed vacuum state, to generate interesting states, including entangled superpositions of two coherent states (Wang et al., 2016). In future work, it will also be interesting to study the effects of postselected von Neumann measurements on other Gaussian and non-Gaussian multipartite continuous-variable radiation fields (Lvovsky et al., 2020; Ra et al., 2020; Walschaers, 2021), as well as their practical applications and implementations in the laboratory.
@@ Appendix A Related expression
Here, we derive explicit expressions for the relevant average values associated with our proposed state $|\Psi\rangle$. However, due to their complexity, many of these expressions are too cumbersome to include in the main text. Therefore, we have provided them in this appendix for reference.
@@
$P_{12}$ $=\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma}{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma L_{n+\delta}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma}{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma L_{n+\delta-1}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$P_{22}$ $=\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$P_{32}$ $=\mathcal{N}_{n}\mathcal{N}_{n-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n-1+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(0)}(\Gamma^{2})$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{n}\mathcal{N}_{n-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(1)}(\Gamma^{2}).$
@@
$P_{42}$ $=\mathcal{N}_{n}\mathcal{N}_{n-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n-2+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(0)}\left(\Gamma^{2}\right)$
$-\Gamma^{2}\mathcal{N}_{n}\mathcal{N}_{n-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n-1+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{4}}{4}\mathcal{N}_{n}\mathcal{N}_{n-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(2)}\left(\Gamma^{2}\right).$
@@
$P_{52}$ $=\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta-1)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{11}$ $=\Gamma^{2}\mathcal{N}_{n}\mathcal{N}_{n-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|{}^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(2)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{n}\mathcal{N}_{n-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|{}^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right).$
@@
$I_{21}$ $=\Gamma^{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-2)}^{(2)}\left(\Gamma^{2}\right)$
$+\Gamma^{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{31}$ $=\mathcal{N}_{n}\mathcal{N}_{n-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n+\delta)!}\Gamma^{2}L_{n+\delta-2}^{(2)}\left(\Gamma^{2}\right)e^{-\frac{\Gamma^{2}}{2}}.$
@@
$I_{41}$ $=-\Gamma\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma}{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{51}$ $=\Gamma\mathcal{N}_{n}\mathcal{N}_{n-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right).$
@@
$K_{12}$ $=\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-2)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\Gamma\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta)}^{(1)}\left(\Gamma^{2}\right)+\Gamma\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}(n+\delta-1)e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma^{2}L_{(n+\delta)}^{(2)}\left(\Gamma^{2}\right)+\frac{\Gamma^{2}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma^{2}L_{(n+\delta-2)}^{(2)}\left(\Gamma^{2}\right)$
$+\Gamma^{2}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right)-\frac{\Gamma^{3}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{3}}{4}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)+\frac{\Gamma^{4}}{16}\mathcal{N}_{n}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$

### v2

@@ VIII Discussion
(4) Weak coupling strength parameter ($\Gamma\ll 1$) with small state parameter $\gamma$ ($|\gamma|<1$). Here, as both the coupling parameter $\Gamma<1$ and $\gamma$ are very small, we can neglect $\Gamma\gamma^{1+\delta/2}$ and higher-order terms in our system state. In this extreme case, the expression for our output state $|\Psi\rangle$ further changes to (unnormalized):
@@
In this regime, the WVA is dominant, and $\Gamma Re[\langle\sigma_{x}\rangle_{w}]$ can be take big quantity compared to $\gamma$ with low success probability of postselection, especially if the pre and post-selected system states are nearly orthogonal. If we assume the PND equal zero ($\delta=0$) and , the $a$-mode is prepared in a superposition state of $|0\rangle_{a}$ and $|1\rangle_{a}$, while the $b$-mode remains in the vacuum state:
$|\phi^{\prime\prime\prime}\rangle_{a}\thicksim\langle 0|_{b}\left(|\Psi^{\prime\prime\prime}\rangle\right)\thicksim\left|0\right\rangle_{a}+\frac{\Gamma}{2}Re[\langle\sigma_{x}\rangle_{w}]\left|1\right\rangle_{a}.$ (53)
@@
Furthermore, under sufficiently large anomalous weak value, it is possible to detect a single photon from the $a$-mode. This feature could be useful in quantum computation processes, as discussed in (Kok et al., 2007). More interestingly, in this extreme case, we can also prepare the state $\left|1+\delta\right\rangle_{a}\left|1\right\rangle_{b}$, but with a lower success probability of postselection. For this extreme case, the advantages of our scheme are highlighted by the very small value regimes shown in the numerical results of Figs. 6 (a) and 6 (b). As we observed, the above prepared states all strongly depends on PND $\delta$ of PCS, and we can obtain various displaced photon numbers states and their superposition states just by changing $\delta$ without photon addition or photon subtraction processes.
In recent study (Chuong and Duc, 2023), the enhancement of non-Gaussianity and nonclassicality of PCS is investigated by adding $k$ photons on $a$-mode while subtracting $l$ photons from the b-mode, and showed that the features including squeezing, quantum correlations and phase space distributions of initial PCS are improved if $k$ and $l$ both greater than one (i.e. $k,l\gg 1$). Furthermore, in Ref. (Truong and Tran, 2024) the same task is considered via subtraction photons from two modes of PCS simultaneously and obtained similar results as in (Chuong and Duc, 2023). However, it’s important to mention that photon addition and/or subtraction operation are probabilistic processes, generally having low success rates even for single photon addition and/or subtraction case. For example, the non-Gaussian optical states can be generated via heralding schemes using photon detectors, with typically use beam splitters to realize the photon subtraction (Horiuchi, 2017). In this case, the transmissivity parameters must be considered when modeling the photon subtraction through a beam splitter, which dramatically reduces the probability of successful state optimization. The accomplishments of above mentioned multiphoton addition and/or subtraction processes need more than one beam splitters and higher precision to photon detectors. Needless to say, such multiphoton addition and/or subtraction operation based optimization to the PCS have very low success efficiency.
Compared with previous state optimization methods, our postselected von Neumann measurement based optimization on PCS is easy to preform and doesn’t need any photon addition and/or subtraction operations. In our scheme the output state detected probability is not very low [see Fig.13] and the efficiency also can be controlled by adjusting the coupling strength parameter and weak values. However, our state optimization method also has its limitations. With the existence of postselection process, we can select partial information based on our wanted outcome and discard the information from the initially prepared state. On the other hand, the large anomalous weak values are often accompanied by a low probability of successful postselection. Due to these limitations, obtaining the desired results via this method requires a large number of repeated measurements on the prepared systems as any precision measurements accomplished with postselected weak measurement (Dixon et al., 2009; Wang et al., 2016b).
@@
We also note that our theoretical scheme could be realized on an optical platform, as the initial system preparation, post-selection of the measured system, and weak coupling between the measured system and MD can all be accomplished in optical labs, as in previous WVA experiments (Xu et al., 2013; Zhu et al., 2021). Furthermore, it may also be possible to implement our scheme in a trapped ion system, as the von Neumann-type interaction Hamiltonian $H_{int}=g\sigma_{x}\otimes P_{x}$ can be prepared in two-dimensional harmonic oscillators of trapped ion systems (Wu et al., 2019; Pan et al., 2020; Su et al., 2023; Jeon et al., 2024; Wu et al., 2024).
@@
Finally, the postselected von Neumann measurement utilizes the technique of WVA, which enables the PCS to exhibit excellent non-classical properties. The PCS under postselected von Neumann measurement can show improved performance in squeezing, quantum statistics, and entanglement by selecting appropriate parameters. This method may provide a valuable quantum resource for checking and enhancing the efficiency of associated quantum tasks based on PCS (Tara and Agarwal, 1994; GÁBRIS and AGARWAL, 2007; Albert et al., 2019).
@@
In this paper, we have proposed a theoretical model to enhance the non-Gaussianity and nonclassicality of PCS using postselected von Neumann measurement. We apply a postselected measurement to one mode (the $a$-mode) of the PCS and investigate related properties, including squeezing, quantum statistics, and entanglement. The results show that after the postselected von Neumann measurement, the associated characteristics of the measurement output state are enhanced in the WM regime with anomalous weak values. Quadrature squeezing achieves better-squeezing effects than the PCS over a wide range of state parameters $\gamma$ under WM. The sum squeezing effect occurs due to WVA for larger weak values. Numerical results using HZ correlation and EPR correlation also show that the entanglement between the two modes of the postselected measurement-enhanced PCS becomes stronger than the initial state. We confirm these findings by examining the scaled joint Wigner function for different system parameters. We also analyze the changes in the initial state after the measurement process in terms of fidelity and observe that as the coupling strength parameter $\Gamma$ increases, the initial state $|\phi\rangle$ changes and can even become orthogonal to $|\phi\rangle$. From this perspective, we also discuss the potential applications of our scheme in quantum state preparation processes and found that various quantum states with high probability could be generated with some additional measurement procedures. Interestingly, we also noticed that with our enhanced PCS the photon addition non-Gaussian states also can be generated without actual photon addition operations.
Furthermore, as an application example of the PCS under postselected von Neumann measurement, we also explore the quantum teleportation protocol by using the postselected measurement-enhanced PCS as quantum channel. The results showed that in weak measurement regimes the successful teleportation of a coherent state via our enhanced PCS-based quantum channel for anomalous weak values could be implemented, but the corresponding averages fidelity still lower than the initial PCS case. Even though this drawback, the existence of the postselected von Neumann measurement could lead the average fidelity be higher than the minimum limit value for successful quantum teleportation, so that the teleportation of the associated quantum information with our enhanced PCS-based quantum channel is still guaranteed.
We explored the possibility of performing postselected von Neumann measurements in multimode radiation fields. For simplicity, we only considered the postselected measurement on one mode of the PCS. It would be interesting to extend this to postselected measurements on both modes of two-mode states, such as the two-mode squeezed vacuum state, to generate interesting states, including entangled superpositions of two coherent states (Wang et al., 2016a). In future work, it will also be interesting to study the effects of postselected von Neumann measurements on other Gaussian and non-Gaussian multipartite continuous-variable radiation fields (Lvovsky et al., 2020; Ra et al., 2020; Walschaers, 2021), as well as their practical applications and implementations in the laboratory.
@@ Appendix A Related expression
Here, we derive explicit expressions for the relevant average values and teleportation fidelity $F_{av}$ associated with our proposed state $|\Psi\rangle$. However, due to their complexity, many of these expressions are too cumbersome to include in the main text. Therefore, we have provided them in this appendix for reference.
@@
$P_{12}$ $=\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma}{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma L_{n+\delta}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma}{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma L_{n+\delta-1}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$P_{22}$ $=\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$P_{32}$ $=\mathcal{N}_{\delta}\mathcal{N}_{\delta-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n-1+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(0)}(\Gamma^{2})$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{\delta}\mathcal{N}_{\delta-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(1)}(\Gamma^{2}).$
@@
$P_{42}$ $=\mathcal{N}_{\delta}\mathcal{N}_{\delta-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n-2+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(0)}\left(\Gamma^{2}\right)$
$-\Gamma^{2}\mathcal{N}_{\delta}\mathcal{N}_{\delta-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n-1+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{4}}{4}\mathcal{N}_{\delta}\mathcal{N}_{\delta-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-2}^{(2)}\left(\Gamma^{2}\right).$
@@
$P_{52}$ $=\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta-1)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta-1}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{11}$ $=\Gamma^{2}\mathcal{N}_{\delta}\mathcal{N}_{\delta-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|{}^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(2)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma^{2}}{2}\mathcal{N}_{\delta}\mathcal{N}_{\delta-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|{}^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right).$
@@
$I_{21}$ $=\Gamma^{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-2)}^{(2)}\left(\Gamma^{2}\right)$
$+\Gamma^{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{31}$ $=\mathcal{N}_{\delta}\mathcal{N}_{\delta-2}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-2}|\gamma|^{2n+\delta}}{(n-2)!(n+\delta)!}\Gamma^{2}L_{n+\delta-2}^{(2)}\left(\Gamma^{2}\right)e^{-\frac{\Gamma^{2}}{2}}.$
@@
$I_{41}$ $=-\Gamma\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$-\frac{\Gamma}{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$
@@
$I_{51}$ $=\Gamma\mathcal{N}_{\delta}\mathcal{N}_{\delta-1}\sum\limits_{n=0}^{\infty}\frac{(\gamma^{*})^{-1}|\gamma|^{2n+\delta}}{(n-1)!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right).$
@@
$K_{12}$ $=\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-2)!}e^{-\frac{\Gamma^{2}}{2}}L_{n+\delta}^{(0)}\left(\Gamma^{2}\right)$
$-\Gamma\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta)}^{(1)}\left(\Gamma^{2}\right)+\Gamma\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}(n+\delta-1)e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{2}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{-\frac{\Gamma^{2}}{2}}\Gamma^{2}L_{(n+\delta)}^{(2)}\left(\Gamma^{2}\right)+\frac{\Gamma^{2}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma^{2}L_{(n+\delta-2)}^{(2)}\left(\Gamma^{2}\right)$
$+\Gamma^{2}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta-1)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right)-\frac{\Gamma^{3}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta)}^{(1)}\left(\Gamma^{2}\right)$
$+\frac{\Gamma^{3}}{4}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}\Gamma L_{(n+\delta-1)}^{(1)}\left(\Gamma^{2}\right)+\frac{\Gamma^{4}}{16}\mathcal{N}_{\delta}^{2}\sum\limits_{n=0}^{\infty}\frac{|\gamma|^{2n+\delta}}{n!(n+\delta)!}e^{\frac{-\Gamma^{2}}{2}}L_{(n+\delta)}^{(0)}\left(\Gamma^{2}\right).$